OpenStudy (anonymous):

Prove that cot(pi) is undefined using an identity. Which identity did you use?

OpenStudy (jdoe0001):

recall that \(\bf cot(\theta ) = \cfrac{cos(\theta)}{sin(\theta)}\)

OpenStudy (anonymous):

But then I have to prove cot(pi) is undefined with an identity. How should I do that?

OpenStudy (jdoe0001):

well.... what's the value of \(\bf sin(\pi) \quad ? \)

OpenStudy (anonymous):

0. So then the denominator would be 0 so then its undefined. Thanks!

OpenStudy (jdoe0001):

yw